# Bridging Continuous and Discrete Tokens for Autoregressive Visual Generation

### v2

@@ 2 Related Work
Concurrent to our work, Fractal Generative Models [18] also employ nested autoregressive structures for image generation. This approach recursively divides the generation process across spatial scales to handle pixel-by-pixel generation, while our method applies dimension-wise autoregressive prediction to model the large token space from post-training quantization. Both methods leverage hierarchical AR modeling for efficiency but focus on different aspects: spatial decomposition versus feature dimension factorization.
@@ 3.1 Post-Training Quantization
Quantization. For a feature vector ${\bm{x}}=\big{(}x^{1},x^{2},\cdots,x^{C}\big{)}$ from the feature map ${\bm{X}}$, we first normalize each dimension to facilitate quantization based on Gaussian properties. We identified practical bounds $[\alpha_{\text{min}},\alpha_{\text{max}}]$ from our experiments, then map these to $[-r,r]$ (where $r=3$ corresponds to three standard deviations):
@@
For each interval $[b_{i},b_{i+1}]$, we compute a reconstruction value as the expected value within that range:
@@ 4.1 Implementation Details
Tokenizer. For VAE, we use the KL-regularized tokenizer from LDM [31] with pre-trained weights from [17]. The tokenizer maps 256×256 images to 16×16 tokens with 16-dimensional feature vectors. For dimension-wise quantization, we set $[\alpha_{\text{min}},\alpha_{\text{max}}]=[-5,5]$, $r=3$, and $B=64$ based on reconstruction quality experiments in Sec. 4.2.
@@ 4.2 Properties of Our Tokenizer
Comparison of Continuous and Discrete Tokenizers. In Fig. 5, we compare the reconstruction quality of typical continuous and discrete tokenizers using ImageNet [6] validation samples. For continuous tokenization, we use the VAE from [17]. For discrete tokenizers, we use VQGAN from LlamaGen’s [36] implementation (an improved version of the original [7]), along with LFQ from Open-MAGVIT2 [20], as the original MAGVIT2 [49] is not open-sourced. As shown in the figure, continuous tokenizers preserve more details, particularly in text and facial features, where discrete tokenizers often struggle. However, our discrete tokenizer achieves reconstruction quality comparable to its continuous counterpart.
@@
(a) Different quantization levels. Finer quantization granularity improves performance, with $B$=64 matching continuous VAE.
@@
Ours ±$\pm{2}$ 2.26 267.9
Ours ±$\pm{3}$ 1.22 299.7
Ours ±$\pm{4}$ 1.13 305.3
Ours ±$\pm{5}$ 1.11 305.4
(b) Different quantization ranges. Feature range ±$\pm 5$ achieves reconstruction quality matching continuous VAE.
Table 1: Ablations on post-training quantization.
Parameters for Tokenizer Quantization. Tab. 1 (a) and Fig. 6 demonstrate the impact of quantization levels $B$(described in Eq. 2 ) on reconstruction quality. One observation is that global structure remains well-preserved across all quantization levels, with differences primarily in fine detail preservation. With $B=8$, significant information loss occurs (rFID=3.69), visible as artifacts in textures and edges. At $B=16$, quality improves substantially (rFID=1.33), with only minor detail loss visible while only 0.2 rFID higher than the continuous baseline. Reconstructions at $B=32$ and $B=64$ become visually indistinguishable from the original inputs, with $B=64$ (rFID=1.11) matching the continuous VAE baseline perfectly. These results show that our post-training quantization approach can achieve continuous-quality reconstruction with enough quantization levels.
Tab. 2 (b) examines how feature range $[\alpha_{\min},\alpha_{\max}]$ affects quantization. The approximately Gaussian distribution of VAE features (due to KL regularization) enables defining a finite range. A narrow range of ±$\pm{2}$ results in information loss (rFID=2.26), while expanding to ±$\pm{5}$ allows our method to match the continuous baseline (rFID=1.11).
@@ 4.3 Properties of Our Generator
training-quantized discrete tokens VQ Mask. CE MaskGIT [3] 177M 6.18 182.1 - -
VQ AR CE RQTran [16] 3.8B 7.55 134.0 - -
VQ AR CE ViT-VQGAN [47] 1.7B 4.17 175.1 - -
VQ Mask. CE TiTok-128 [50] 287M 1.97 281.8 - -
LFQ AR CE MAGVIT-v2 [49] 307M 1.78 319.4 - -
LFQ AR CE Open-MAGVIT2-L [20] 804M 2.51 271.7 0.84 0.54
VQ AR CE LlamaGen [36] 3.1B 2.18 263.3 0.81 0.58
VQ AR CE VAR [38] 2.0B 1.73 350.2 0.82 0.60
continuous-valued tokens VAE Diff. Diff. LDM-4 [31] 400M 3.60 247.7 0.87 0.48
VAE Diff. Diff. U-ViT-H/2-G [1] 501M 2.29 263.9 0.82 0.57
VAE Diff. Diff. DiT-XL/2 [24] 675M 2.27 278.2 0.83 0.57
VAE Diff. Diff. MDTv2-XL/2 [8] 676M 1.58 314.7 0.79 0.65
VAE AR GMM GIVT [39] 304M 3.35 - 0.84 0.53
VAE AR Flow FlowAR-H [29] 1.9B 1.65 296.5 0.83 0.60
VAE AR Diff. MAR-L [17] 479M 1.78 296.0 0.81 0.60
VAE AR Diff. MAR-H [17] 943M 1.55 303.7 0.81 0.62
post-training quantized discrete tokens VAE AR CE Ours-L 486M 1.76 294.8 0.80 0.63
VAE AR CE Ours-H 910M 1.55 313.3 0.80 0.65
@@
Channel Grouping. We explore whether combining multiple channels for joint classification could improve generation performance. Due to the exponential growth in class count when grouping channels, we conduct experiments with $B$=16 quantization on a 172M parameter backbone. As shown in Tab. 2(d), our default approach performs classification for each channel. Grouping channels exponentially increases vocabulary size: 2 channels/group requires predicting from 256 classes, while 4 channels/group expands to 65,536 classes. Despite increased parameters (530M vs. 60M for single-channel), generation quality worsens (gFID from 2.28 to 3.24). We suspect this is because larger classification spaces pose greater modeling challenges for the generator, despite the increased parameter count. These results validate the effectiveness of our dimension-wise factorization approach for handling large vocabulary spaces.
@@ Acknowledgment
The authors are grateful to Tianhong Li for helpful discussions on MAR and to Yi Jiang, Difan Zou, and Yujin Han for valuable feedback on the early version of this work.
@@
  • Sec. C discusses limitations and broader impacts.
  • Sec. D showcases additional image generation results.
@@ B Speed Comparison of Token prediction
C Limitations and Broader Impacts
@@
D More Visualization Results

### v3

@@ 3.1 Post-Training Quantization
Quantization. For a feature vector ${\bm{x}}=\big{(}x^{1},x^{2},\cdots,x^{C}\big{)}$ from the feature map $\bm{X}$}, we first normalize each dimension to facilitate quantization based on Gaussian properties. We identified practical bounds $\alpha_{\text{min}},\alpha_{\text{max}}]$ from our experiments, then map these to [r,r][-r,r] (where $r=3$ corresponds to three standard deviations):
@@
For each interval $b_{i},b_{i+1}]$, we compute a reconstruction value as the expected value within that range:
@@ 4.1 Implementation Details
Tokenizer. For VAE, we use the KL-regularized tokenizer from LDM [31] with pre-trained weights from [18]. The tokenizer maps 256×256 images to 16×16 tokens with 16-dimensional feature vectors. For dimension-wise quantization, we set $\alpha_{\text{min}},\alpha_{\text{max}}]=[-5,5]$, $r=3$, and $B=64$ based on reconstruction quality experiments in Sec. 4.2.
@@ 4.2 Properties of Our Tokenizer
Comparison of Continuous and Discrete Tokenizers. In Fig. 5, we compare the reconstruction quality of typical continuous and discrete tokenizers using ImageNet [7] images. For continuous tokenization, we use the VAE from [18]. For discrete tokenizers, we use VQGAN from LlamaGen’s [36] implementation (an improved version of the original [8]), along with LFQ from Open-MAGVIT2 [20], as the original MAGVIT2 [50] is not open-sourced. As shown in the figure, continuous tokenizers preserve more details, particularly in text and facial features, where discrete tokenizers often struggle. However, our discrete tokenizer achieves reconstruction quality comparable to its continuous counterpart.
@@
(a) Different quantization levels. Finer quantization granularity improves performance, with BB=64 matching continuous VAE.
@@
Ours $\pm{2}$ 2.26 267.9
Ours $\pm{3}$ 1.22 299.7
Ours $\pm{4}$ 1.13 305.3
Ours $\pm{5}$ 1.11 305.4
(b) Different quantization ranges. Feature range $\pm 5$ achieves reconstruction quality matching continuous VAE.
Table 1: Ablations on post-training quantization on the ImageNet Training Set.
Parameters for Tokenizer Quantization. Tab. 1 (a) and Fig. 6 demonstrate the impact of quantization levels BB(described in Eq. 2 ) on reconstruction quality. One observation is that global structure remains well-preserved across all quantization levels, with differences primarily in fine detail preservation. With $B=8$, significant information loss occurs (rFID=3.69), visible as artifacts in textures and edges. At $B=16$, quality improves substantially (rFID=1.33), with only minor detail loss visible while only 0.2 rFID higher than the continuous baseline. Reconstructions at $B=32$ and $B=64$ become visually indistinguishable from the original inputs, with $B=64$ (rFID=1.11) matching the continuous VAE baseline perfectly. These results show that our post-training quantization approach can achieve continuous-quality reconstruction with enough quantization levels.
Tab. 2 (b) examines how feature range $\alpha_{\min},\alpha_{\max}]$ affects quantization. The approximately Gaussian distribution of VAE features (due to KL regularization) enables defining a finite range. A narrow range of $\pm{2}$ results in information loss (rFID=2.26), while expanding to $\pm{5}$ allows our method to match the continuous baseline (rFID=1.11). Note that to align with MAR [18], the above metrics are evaluated on the training set of ImageNet. For comparison with other works, our best rFID on the validation set is 0.53.
@@ 4.3 Properties of Our Generator
Channel Grouping. We explore whether combining multiple channels for joint classification could improve generation performance. Due to the exponential growth in class count when grouping channels, we conduct experiments with BB=16 quantization on a 172M parameter backbone. As shown in Tab. 2(d), our default approach performs classification for each channel. Grouping channels exponentially increases vocabulary size: 2 channels/group requires predicting from 256 classes, while 4 channels/group expands to 65,536 classes. Despite increased parameters (530M vs. 60M for single-channel), generation quality worsens (gFID from 2.28 to 3.24). We suspect this is because larger classification spaces pose greater modeling challenges for the generator, despite the increased parameter count. These results validate the effectiveness of our dimension-wise factorization approach for handling large vocabulary spaces.
@@ Acknowledgment
This work is partially supported by the National Nature Science Foundation of China (No. 62402406). The authors are grateful to Tianhong Li for helpful discussions on MAR and to Yi Jiang, Difan Zou, and Yujin Han for valuable feedback on the early version of this work.
@@
  • Sec. C evaluates generalization to different VAE and AE architectures.
  • Sec. D discusses limitations and broader impacts.
  • Sec. E showcases additional image generation results.
@@ B Speed Comparison of Token prediction
C Generalization to Different VAE and AE Architectures
To evaluate the generalization of our post-training quantization approach, we select two representative alternative autoencoders for evaluation: VAVAE [47], a state-of-the-art VAE with representation alignment, and DCAE [4], achieving high compression rates without KL loss constraints. Fig. 9 visualizes the latent feature distributions of different autoencoders. Although the value ranges differ across architectures, all exhibit similar near-Gaussian distributions. This consistency validates that the bounded, approximately Gaussian property, independent of specific architectural designs or training constraints like KL regularization. As described in the Method section, even linear quantization achieves good reconstruction results, demonstrating the robustness of our approach across different quantization schemes.
Tab. 6 shows reconstruction results after applying our quantization with corresponding rescaling and quantization granularity. Our method successfully preserves reconstruction quality across different architectures: VAVAE achieves identical performance (rFID=0.28) to its continuous baseline using B=128 and r=3.5, while DCAE matches its baseline (rFID=0.77) with B=64 and r=8. These results demonstrate that our post-training quantization approach generalizes effectively across diverse autoencoder architectures while maintaining reconstruction fidelity.
Figure 9: Latent distributions of different autoencoders. Despite architectural differences and training objectives, all three models exhibit similar near-Gaussian distributions, validating the generalizability of our quantization approach.
AE Ori. FID B Range TokenBridge FID
VAVAE 0.28 128 [-3.5, 3.5] 0.28
DCAE 0.77 64 [-8, 8] 0.77
Table 6: Reconstruction quality across different autoencoder architectures. Our post-training quantization preserves reconstruction fidelity when applied with appropriate parameters.
D Limitations and Broader Impacts
@@
E More Visualization Results